Complete all the valence bond hybridization and bonding scheme for the following:

a. H2CO

b.HCCH

c.HONO

d.BRF4-

e.NO2+
Answer #1

Sol :-

Hybridization (H) = Number of lone pairs around central atom + Number of bond pairs around central atom.

# If H = 2 , then hybridization is sp and molecular shape is linear.

# If H = 3 , then hybridization is sp2 and molecular shape is trigonal planar.

# If H = 4 , then hybridization is sp3 and molecular shape is tetrahedral.

# If H = 5 , then hybridization is sp3d and molecular shape is trigonal bipyramidal.

# If H = 6 , then hybridization is sp3d2 and molecular shape is octahedral

# If H = 7 , then hybridization is sp3d3 and molecular shape is pentagonal pyramidal.

(a). In H2CO :

Hybridization (H) = 0 + 3 = 3, then hybridization of central atom carbon is sp2.

and

Molecuar shape is = Trigonal planar.

----------------------------------------

(b). In HCCH :

Hybridization (H) = 0 + 2 = 2, then hybridization of central atom carbon is sp.

and

Molecuar shape is = Linear.

----------------------------------------

(c). In HONO :

Hybridization (H) = 1 + 2 = 2, then hybridization of central atom carbon is sp2.

and

Molecuar shape is = Bent.

----------------------------------------

(d). In BrF4- :

Hybridization (H) = 2 + 4 = 6, then hybridization of central atom carbon is sp3d2.

and

Molecuar shape is = Sequare planar.

----------------------------------------

(e). In NO2+ :

Hybridization (H) = 0 + 2 = 2, then hybridization of central atom carbon is sp.

and

Molecuar shape is = Linear.

----------------------------------------

H-CE C-H =öz OH (C:) Từ 0 1 :